OpenStudy (anonymous):

A rectangular shaped flower garden has a length of 5 yards and an area of 15 square yards. A neighbors flower garden is similar and has a length of 7 yards. What is the area of the neighbors flower garden? round your answer to the nearest whole number.

OpenStudy (anonymous):

What do you mean by similar? Do you mean it has the same area?

OpenStudy (cwrw238):

or same shape?

OpenStudy (anonymous):

Im guessing this what i needed help on. :/

OpenStudy (anonymous):

Well, if it was size then.. 35 yd squared. it's more complicated with same area

OpenStudy (anonymous):

Thank you

OpenStudy (cwrw238):

if its same shape then areas are in ratio 5^2:7^2 or 25 : 49 ie 15/x = 25/49 x = 49*15/25 = 49*3 / 5 = 147/5 = 29.4 sq yds